OpenStudy (anonymous):

The atomic masses of 79 Br and 81 Br are 78.918336 and 80.916290 amu, respectively. The average atomic mass of bromine is 79.904 amu. Calculate the natural abundances of these two isotopes.

OpenStudy (anonymous):

Ok so the formula is: mass of Br-79 amu times % natural abundance of Br-79 + mass of Br-81 amu times % natural abundance of Br-81=total amu so since you know the masses you can put those in, but you don't know the natural abundance but, using algebra you can use a letter coefficient and create the equation: \[(78.918336 \times x) + (80.916290 \times \left( 100-x \right)) = 79.904\] and solve for x

OpenStudy (anonymous):

does that help?